OpenStudy (anonymous):

MEDALS + FAN!!!!!!!!!!!!! Can ANY1 HELP ME WITH MY MIDDLE SCHOOL HOME WORK PLEASE JUST 2 OR 1 QUESTIONS????

OpenStudy (anonymous):

yeah...

OpenStudy (anonymous):

the last 1

OpenStudy (anonymous):

First, you would turn the fractions into mixed numbers... You would get: 13/1 for 6 1/2 and 13/4 for 3 1/4... Then you would divide those fractions using the Keep, Switch, Flip method... Keep the 13/1... Switch the division symbol to a multiplication symbol... And Flip the fraction (the reciprocal): 3 1/4 = 4/13... And then you will multiply... 13/1 * 4/13 Which will equal... 4/1 = 4 So 4 is the answer. :)

OpenStudy (anonymous):

can u help me with more? @gold_grace123

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

@gold_grace123

OpenStudy (anonymous):

which one

OpenStudy (anonymous):

9 10 11 @gold_grace123

OpenStudy (anonymous):

9~ 7/10 10~ D 11~ G

OpenStudy (anonymous):

the rest @gold_grace123

OpenStudy (anonymous):

b

OpenStudy (anonymous):

i think